OpenStudy (anonymous):

Is this function right?? Let f(x)=9x-7. Simplify the expression f(x+h) -f(x)/h 9(x+h)-7[9x-7]/h = 9 Is the answer 9??

OpenStudy (anonymous):

What is the question? You asked like 3..

OpenStudy (anonymous):

this is the question..Let f(x)=9x-7. Simplify the expression f(x+h) -f(x)/h that was me showing work

OpenStudy (akashdeepdeb):

f(x) = 9x - 7 f(x+h) = 9(x+h) - 7 We have to find: \(\frac{f(x+h) - f(x)}{h}\) \[= \frac{9(x+h) - 7 - 9x + 7}{h} = \frac{9x + 9h - 7 - 9x + 7}{h} = \frac{9h}{h} = 9\] 9 is the first differential of f(x)

OpenStudy (anonymous):

so it's 9 right @AkashdeepDeb

OpenStudy (akashdeepdeb):

Yes.